OpenStudy (anonymous):

Wat is the slope of a line that passes throught the point (-1,1) and is parallel to a line that passes throught (4,6) and (-1,-4)?

OpenStudy (anonymous):

Someone plz help!

OpenStudy (anonymous):

10/5=2/1=2 slope is 2

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

No Problem(=

OpenStudy (anonymous):

Shouldn't it be -1/2?

OpenStudy (anonymous):

Nevermind, I read that as perpendicular.